OpenStudy (unofficialllyy):

A textbook store sold a combined total of 434 psychology and biology textbooks in a week. The number of biology textbooks sold was 86 less than the number of psychology textbooks sold. How many textbooks of each type were sold?

OpenStudy (unofficialllyy):

@jim_thompson5910

OpenStudy (unofficialllyy):

@texaschic101

jimthompson5910 (jim_thompson5910):

`A textbook store sold a combined total of 434 psychology and biology textbooks in a week.` let x = # of psychology books y = # of biology books what equation can we form from the information given above?

OpenStudy (unofficialllyy):

x+y=434?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

`The number of biology textbooks sold was 86 less than the number of psychology textbooks sold.` what equation can we form here?

OpenStudy (unofficialllyy):

x+y-86=434?

jimthompson5910 (jim_thompson5910):

`The number of biology textbooks sold was 86 less than the number of psychology textbooks sold` x = # of psychology books y = # of biology books (number of bio books) = (number of psych books) - 86 (y) = (x) - 86 y = x - 86 so we can form the equation y = x-86 agreed?

OpenStudy (unofficialllyy):

yes

jimthompson5910 (jim_thompson5910):

the system of equations is then x+y = 434 y = x-86

jimthompson5910 (jim_thompson5910):

\[\Large \color{red}{y} =\color{red}{x-86}\] ------------------------ \[\Large x+y = 434\] \[\Large x+\color{red}{y} = 434\] \[\Large x+\color{red}{x-86} = 434\] \[\Large 2x-86 = 434\] \[\Large x = ???\]

OpenStudy (raven7266):

x=260

OpenStudy (unofficialllyy):

thank you @jim_thompson5910